# Transport, Variational Inference and Diffusions: with Applications to Annealed Flows and Schrödinger Bridges

### v1

@@ 2.2 SDE transition densities – hierarchical VAEs in the infinite depth limit
In this section, we take inspiration from Section 2.1 and Tzen and Raginsky (2019); Li et al. (2020); Huang et al. (2021a) to develop the infinite-depth limit ($L\rightarrow\infty$), using the machinery of stochastic differential equations (SDEs). To this end, we think of $l=0,\ldots,L$ as discrete instances in a fixed time interval $[0,T]$, equidistant with time step $\delta$, that is, we set $\delta=TL^{-1}$. The discrete paths ${\bm{y}}_{0:L}$ give rise to continuous paths $({\bm{Y}}_{t})_{0\leq t\leq T}\in C([0,T];\mathbb{R}^{d})$ by setting ${\bm{Y}}_{\delta l}={\bm{y}}_{l}$ and interpolating linearly between ${\bm{Y}}_{\delta l}$ and ${\bm{Y}}_{\delta(l+1)}$. To complete the set-up, we think of $a^{\phi}=(a^{\phi}_{0},\ldots,a^{\phi}_{L-1})$ and $b^{\theta}=(b_{1}^{\theta},\ldots,b^{\theta}_{L})$ in (2.1) as arising from time-dependent vector fields $a,b\in C^{\infty}([0,T]\times\mathbb{R}^{d};\mathbb{R}^{d})$ via $a^{\phi}_{l}({\bm{y}}_{l})=a_{t\delta^{-1}}({\bm{Y}}_{\delta l})$ and $b^{\theta}_{l}({\bm{y}}_{l})=b_{t\delta^{-1}}({\bm{Y}}_{\delta l})$.
@@
We are now prepared to invoke the limit $\delta\rightarrow 0$ (equivalently, $L\rightarrow\infty$), keeping $T>0$ fixed. The proposed relabelling and rescaling transforms the Markov chains in (2.1) into continuous-time dynamics described by the SDEs
${\mathrm{d}}{\bm{Y}}_{t}$ $=a_{t}({\bm{Y}}_{t})\,{\mathrm{d}}t+\sigma\overrightarrow{\mathrm{d}}{\bm{W}}_{t},\quad{\bm{Y}}_{0}\sim\mu$
$\implies({\bm{Y}}_{t})_{0\leq t\leq T}\sim\mathbb{Q}^{\mu,a}\equiv\overrightarrow{\mathbb{P}}^{\mu,a},$ (13a)
${\mathrm{d}}{\bm{Y}}_{t}$ $=b_{t}({\bm{Y}}_{t})\,{\mathrm{d}}t+\sigma\overleftarrow{\mathrm{d}}{\bm{W}}_{t},\quad{\bm{Y}}_{T}\sim\nu$
$\implies({\bm{Y}}_{t})_{0\leq t\leq T}\sim\mathbb{P}^{\nu,b}\equiv\overleftarrow{\mathbb{P}}^{\nu,b},$ (13b)
@@
where $\overrightarrow{\mathrm{d}}$ and $\overleftarrow{\mathrm{d}}$ denote forward and backward Itô integration (see Appendix A for more details and remarks on the notation), and $({\bm{W}}_{t})_{0\leq t\leq T}$ is a standard Brownian motion. In complete analogy with (2.1), the SDEs in (2.2) induce the distributions $\mathbb{Q}^{\mu,a}$ and $\mathbb{P}^{\nu,b}$ on the path space $C([0,T];\mathbb{R}^{d})$. Relating back to the discussion in the introduction, note that we maintain the relations ${\bm{Y}}_{0}={\bm{x}}$ and ${\bm{Y}}_{T}={\bm{z}}$, and the transitions are parameterised by the vector fields $a,b$, in the sense that $p^{\theta}({\bm{x}}|{\bm{z}})={\mathbb{P}}^{\nu,b^{\theta}}_{0}({\bm{x}}|{\bm{Y}}_{T}={\bm{z}})={{\mathbb{P}}}^{\delta_{{\bm{z}}},b^{\theta}}_{0}({\bm{x}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})={\mathbb{Q}}^{\mu,a^{\phi}}_{T}({\bm{z}}|{\bm{Y}}_{0}={\bm{x}})={{\mathbb{Q}}}^{\delta_{{\bm{x}}},a^{\phi}}_{T}({\bm{z}})$.
@@
The following well-known result (Anderson, 1982; Haussmann and Pardoux, 1985; Nelson, 1967) allows us to relate forward and backward path measures via a local (score-matching) condition for the reversal relation in (4).  [^2] The global condition $\overrightarrow{{\mathbb{P}}}^{\mu,a}=\overleftarrow{{\mathbb{P}}}^{\nu,b}$ can be captured by the local condition (14) due to the Markovian nature of (2.2).
@@
For fixed $\mu$ and $a$, denote the time-marginals of the corresponding path measure by ${\mathbb{P}}^{\mu,a}_{t}=:\rho^{\mu,a}_{t}$. Assuming that for all $t\in(0,T]$, $\rho_{t}$ admits a strictly positive and smooth density with respect to the Lebesgue measure, we have that $\overrightarrow{{\mathbb{P}}}^{\mu,a}=\overleftarrow{{\mathbb{P}}}^{\nu,b}$ if and only if $\nu=\overrightarrow{{\mathbb{P}}}^{\mu,a}_{T}$ and
$\quad b_{t}=a_{t}-\sigma^{2}\nabla\operatorname{ln}\rho^{\mu,a}_{t},\quad\text{for all }t\in(0,T].$ (14)
@@
A similarly clean characterisation of equality between forward and backward path measures is not available for the discrete-time setting as presented in (2.1). In particular, Gaussianity of the intermediate transitions is not preserved under time-reversal, and thus the parameterisations in (2.1) are sub-optimal without the limit $\delta\rightarrow 0$.
@@
Let $\overrightarrow{{\mathbb{P}}}^{\Gamma_{0},\gamma^{+}}=\overleftarrow{{\mathbb{P}}}^{\Gamma_{T},\gamma^{-}}$ be a reference path measure (that is, $\Gamma_{0}$, $\Gamma_{T}$ and $\gamma^{\pm}$ define diffusions as in (2.2) and are related as in Proposition 2.1), absolutely continuous with respect to both $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ and $\overleftarrow{{\mathbb{P}}}^{\nu,b}$. Then, $\overrightarrow{{\mathbb{P}}}^{\mu,a}$-almost surely, the corresponding Radon-Nikodym derivative (RND) can be expressed as follows,
@@
The proof relies on Girsanov’s theorem (Üstünel and Zakai, 2013), using the reference to relate the forward and backward processes. For details, see Appendix E. ∎
@@
According to Proposition 2.2, the Radon-Nikodym derivative between $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ and $\overleftarrow{{\mathbb{P}}}^{\nu,b}$ can be decomposed into boundary terms involving the marginals in (15a), as well as forward and backward path integrals in (15b) and (15c). Note that the left-hand side of (15a) does not depend on the reference $\Gamma_{0,T}$, $\gamma^{\pm}$. Hence, the expressions in (2.2) are in principle equivalent for all choices of reference, and close inspection shows that the freedom in $\Gamma_{0,T}$ and $\gamma^{\pm}$ allows us to ‘reweight’ between (15a), (15b) and (15c), or even cancel terms. [^3] Note that not all terms can be cancelled at the same time, since $\Gamma_{0,T}$, $\gamma_{\pm}$ cannot be chosen independently (they need to satisfy the conditions imposed by Proposition 2.1). A canonical choice is the Lebesgue measure for $\Gamma_{0}$ and $\Gamma_{T}$, and $\gamma^{\pm}=0$, see Appendix C.1. For constructing algorithms, we may however choose the reference diffusion in such a way that intractable quantities are suppressed, or so that the variance of associated estimators is reduced. We will come back to these points in Section 3.
@@
The distinction between forward and backward integration in (2.2) is related to the time points at which the integrands $\left(a_{t}-\gamma^{+}_{t}\right)({\bm{Y}}_{t})$ and $\left(b_{t}-\gamma^{-}_{t}\right)({\bm{Y}}_{t})$ would be evaluated in discrete-time approximations, e.g.,
$\int_{0}^{T}\!\!a_{t}({\bm{Y}}_{t})\!\cdot\!\overrightarrow{\mathrm{d}}{\bm{Y}}_{t}\approx\sum_{i}a_{t_{i}}({\bm{Y}}_{t_{i}})\!\cdot\!({\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}),$
$\int_{0}^{T}\!\!a_{t}({\bm{Y}}_{t})\!\cdot\!\overleftarrow{\mathrm{d}}{\bm{Y}}_{t}\!\approx\!\sum_{i}a_{t_{i+1}}({\bm{Y}}_{t_{i+1}})\cdot({\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}).$
Alternatively, forward and backward integrals can be transformed into each other using the conversion formula
@@
We refer to Kunita (2019) and Appendix A for further details. In passing, we note that (16) allows us to eliminate the Hutchinson estimator (Hutchinson, 1989) from a variety of objectives, potentially reducing the variance of gradient estimators, see Appendix C.1.
@@
Let us conclude this section by translating Framework 1 into the setting of (2.2), noting that (2.1) continues to hold with appropriate modifications.
Framework ${}^{\prime}$.
@@
For a divergence $D$ on path space, minimise $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})$. If $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})=0$, then (13a) transports $\mu$ to $\nu$, and (13b) transports $\nu$ to $\mu$.
@@
At optimality, $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})=0$, Proposition 2.1 allows us to obtain the scores associated to the learned diffusion via $\sigma^{2}\nabla\operatorname{ln}\rho_{t}^{\mu,a}=a_{t}-b_{t}$. In this way, Framework ${}^{\prime}$ is closely connected to (and in some ways extends) score-matching ideas (Song and Ermon, 2019; Song et al., 2021).
3 Forward-reverse SDEs as a unifying perspective for generative modeling and sampling
We begin by showing that recent approaches towards generative modeling and sampling can be recovered from Framework ${}^{\prime}$ by making specific choices for the divergence $D$, the parameterisations for $a$ and $b$, as well as for the reference diffusion $\overrightarrow{{\mathbb{P}}}^{\Gamma_{0},\gamma^{+}}=\overleftarrow{{\mathbb{P}}}^{\Gamma_{T},\gamma^{-}}$ in Proposition 2.2. Recall from the introduction that complete flexibility in $a$ and $b$ will render the minima of $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})$ highly nonunique, and so it is natural to fix either $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ or $\overleftarrow{{\mathbb{P}}}^{\nu,b}$, or to impose further constraints relating the two (see Section 3.1).
@@ 3 Forward-reverse SDEs as a unifying perspective for generative modeling and sampling
Score-based generative modeling: Letting $\mu$ be the target and fixing the forward drift $a_{t}$, and, motivated by Proposition 2.1, parameterising the backward drift as $b_{t}=a_{t}-s_{t}$, we recover the SGM objectives in Hyvärinen and Dayan (2005); Song and Ermon (2019); Song et al. (2021) from $D=D_{\mathrm{KL}}$; when $\overrightarrow{{\mathbb{P}}}^{\mu,a}=\overleftarrow{{\mathbb{P}}}^{\nu,b}$, the variable drift component $s_{t}$ will represent the score $\sigma^{2}\nabla\operatorname{ln}\rho^{\mu,a}_{t}$. Modifications can be obtained from the conversion formula (16), see Appendix C.2.
@@
Score-based sampling – ergodic drift: In this setting, $\nu$ becomes the target and we fix $b_{t}$ to be the drift of an ergodic (backward) process. Then choosing $\Gamma_{0,T}=\mu$, $\gamma^{\pm}=b$ allows us to recover the approaches in Vargas et al. (2023a); Berner et al. (2022). Possible generalisations based on Framework ${}^{\prime}$ include IWAE-type objectives, see Appendix C.3.
@@
Score-based sampling – Föllmer drift: Finally choosing $b_{t}(x)=x/t$ we can recover the Föllmer drift approach to sampling (Follmer, 1984; Vargas et al., 2023b; Zhang and Chen, 2022; Huang et al., 2021b), see Appendix C.3.
Domain adaptation and stochastic filtering: Setting $\overrightarrow{{\mathbb{P}}}^{\Gamma_{0},\gamma^{+}}$ equal to the signal process in stochastic filtering (Reich and Cotter, 2015), we can target the corresponding Bayesian posterior with log-likelihood $\operatorname{ln}(\nu/\Gamma_{T})$. Note that to obtain a tractable objective, we would need in a preliminary step to learn $\gamma^{-}$ using score matching techniques. Similarly, we can design a reference process with $\Gamma_{T}=\nu$ in case the log-density of $\nu$ is not available, as in domain adaptation (Courty et al., 2017).
3.1 Score-based annealed flows
Departing from the discussion so far, we now attempt to vary both $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ and $\overleftarrow{{\mathbb{P}}}^{\nu,b}$ simultaneously. To retain uniqueness, and guided by Proposition 2.1, we impose the additional constraint $b_{t}=a_{t}-\sigma^{2}\nabla\operatorname{ln}\pi_{t}$, so that when $\overrightarrow{{\mathbb{P}}}^{\mu,a}=\overleftarrow{{\mathbb{P}}}^{\nu,b}$, the time marginals necessarily satisfy $\overrightarrow{{\mathbb{P}}}^{\mu,a}_{t}=\overleftarrow{{\mathbb{P}}}^{\nu,b}_{t}=\pi_{t}$, for all $t\in[0,T]$. In this construction, $(\pi_{t})_{t\in[0,T]}$ is a prescribed curve of distributions whose scores (and unnormalised densities $\hat{\pi}_{t}$) are assumed to be available in tractable form; this is the scenario typically considered in annealed importance sampling and related approaches towards computing posterior expectations (Neal, 2001; Reich, 2011; Heng et al., 2021, 2020; Arbel et al., 2021; Doucet et al., 2022). To complete the setting, we fix the parameterisation $a_{t}=\sigma^{2}\nabla\operatorname{ln}\pi_{t}+\nabla\phi_{t}$ in terms of a scalar function $\phi_{t}$, [^4] Restricting $a_{t}$ further to be of gradient form ensures uniqueness, see Proposition 3.1 below. so that the forward process is given by the controlled time-inhomogeneous Langevin dynamics
${\bm{Y}}_{0}$ $\!\sim\!\pi_{0},$

### v6

@@ 2.2 Diffusion models – hierarchical VAEs in the infinite depth limit
Here we take inspiration from Section 2.1 and Tzen & Raginsky (2019a); Li et al. (2020); Huang et al. (2021a) to investigate the $L\rightarrow\infty$ limit, using stochastic differential equations (SDEs). To this end, we think of $l=0,\ldots,L$ as discrete instances in a fixed time interval $[0,T]$, equidistant with time step $\delta$, that is, we set $\delta=TL^{-1}$. The discrete paths ${\bm{y}}_{0:L}$ give rise to continuous paths $({\bm{Y}}_{t})_{0\leq t\leq T}\in C([0,T];\mathbb{R}^{d})$ by setting ${\bm{Y}}_{\delta l}={\bm{y}}_{l}$ and linearly interpolating ${\bm{Y}}_{\delta l}$ and ${\bm{Y}}_{\delta(l+1)}$. To complete the set-up, we think of $a^{\phi}=(a^{\phi}_{0},\ldots,a^{\phi}_{L-1})$ and $b^{\theta}=(b_{1}^{\theta},\ldots,b^{\theta}_{L})$ in (8) as arising from time-dependent vector fields $a,b\in C^{\infty}([0,T]\times\mathbb{R}^{d};\mathbb{R}^{d})$ via $a^{\phi}_{l}({\bm{y}}_{l})=a_{t\delta^{-1}}({\bm{Y}}_{\delta l})$ and $b^{\theta}_{l}({\bm{y}}_{l})=b_{t\delta^{-1}}({\bm{Y}}_{\delta l})$.
@@
Taking the limit $\delta\rightarrow 0$, while keeping $T>0$ fixed, transforms the Markov chains in (9) into continuous-time dynamics described by the SDEs (Tzen & Raginsky, 2019a)
${\mathrm{d}}{\bm{Y}}_{t}$ $=a_{t}({\bm{Y}}_{t})\,{\mathrm{d}}t+\sigma\overrightarrow{\mathrm{d}}{\bm{W}}_{t},\quad{\bm{Y}}_{0}\sim\mu\implies({\bm{Y}}_{t})_{0\leq t\leq T}\sim\mathbb{Q}^{\mu,a}\equiv\overrightarrow{\mathbb{P}}^{\mu,a},$ (12a)
${\mathrm{d}}{\bm{Y}}_{t}$ $=b_{t}({\bm{Y}}_{t})\,{\mathrm{d}}t+\sigma\overleftarrow{\mathrm{d}}{\bm{W}}_{t},\quad{\bm{Y}}_{T}\sim\nu\implies({\bm{Y}}_{t})_{0\leq t\leq T}\sim\mathbb{P}^{\nu,b}\equiv\overleftarrow{\mathbb{P}}^{\nu,b},$ (12b)
@@
where $\overrightarrow{\mathrm{d}}$ and $\overleftarrow{\mathrm{d}}$ denote forward and backward Itô integration (see Appendix A for more details and remarks on the notation), and $({\bm{W}}_{t})_{0\leq t\leq T}$ is a standard Brownian motion. In complete analogy with (9), the SDEs in (12) induce the distributions $\mathbb{Q}^{\mu,a}$ and $\mathbb{P}^{\nu,b}$ on the path space $C([0,T];\mathbb{R}^{d})$. Relating back to the discussion in the introduction, note that we maintain the relations ${\bm{Y}}_{0}={\bm{x}}$ and ${\bm{Y}}_{T}={\bm{z}}$, and the transitions are parameterised by the vector fields $a,b$, in the sense that $p^{\theta}({\bm{x}}|{\bm{z}})={\mathbb{P}}^{\nu,b^{\theta}}_{0}({\bm{x}}|{\bm{Y}}_{T}={\bm{z}})={{\mathbb{P}}}^{\delta_{{\bm{z}}},b^{\theta}}_{0}({\bm{x}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})={\mathbb{Q}}^{\mu,a^{\phi}}_{T}({\bm{z}}|{\bm{Y}}_{0}={\bm{x}})={{\mathbb{Q}}}^{\delta_{{\bm{x}}},a^{\phi}}_{T}({\bm{z}})$.
@@
The following well-known result (Anderson, 1982; Nelson, 1967) allows us to relate forward and backward path measures via a local (score-matching) condition for the reversal relation in (4).  [^2] The global condition $\overrightarrow{{\mathbb{P}}}^{\mu,a}\!=\!\overleftarrow{{\mathbb{P}}}^{\nu,b}$ is captured by the local condition (13) due to (12)’s Markovian nature.
@@
For $\mu$ and $a$ of sufficient regularity, denote the time-marginals of the corresponding path measure by $\overrightarrow{{\mathbb{P}}}^{\mu,a}_{t}=:\rho^{\mu,a}_{t}$. Then $\overrightarrow{{\mathbb{P}}}^{\mu,a}=\overleftarrow{{\mathbb{P}}}^{\nu,b}$ if and only if
$\nu=\overrightarrow{{\mathbb{P}}}^{\mu,a}_{T}\qquad\text{and}\qquad b_{t}=a_{t}-\sigma^{2}\nabla\operatorname{ln}\rho^{\mu,a}_{t},\qquad\text{for all }t\in(0,T].$ (13)
@@
A similarly clean characterisation of equality between forward and backward path measures is not available for the discrete-time setting as presented in (9). In particular, Gaussianity of the intermediate transitions is not preserved under time-reversal.
@@
Let $\overrightarrow{{\mathbb{P}}}^{\Gamma_{0},\gamma^{+}}=\overleftarrow{{\mathbb{P}}}^{\Gamma_{T},\gamma^{-}}$ be a reference path measure (that is, $\Gamma_{0}$, $\Gamma_{T}$ and $\gamma^{\pm}$ define diffusions as in (12) and are related as in Proposition 2.1), absolutely continuous with respect to both $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ and $\overleftarrow{{\mathbb{P}}}^{\nu,b}$. Then, $\overrightarrow{{\mathbb{P}}}^{\mu,a}$-almost surely, the corresponding Radon-Nikodym derivative (RND) can be expressed as follows,
@@
The proof relies on Girsanov’s theorem (Üstünel & Zakai, 2013), using the reference to relate the forward and backward processes. For details, see Appendix E. ∎
@@
According to Proposition 2.2, the Radon-Nikodym derivative between $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ and $\overleftarrow{{\mathbb{P}}}^{\nu,b}$ can be decomposed into boundary terms (14a), as well as forward and backward path integrals (14b) and (14c). Since the left-hand side of (14a) does not depend on the reference $\Gamma_{0,T}$, $\gamma^{\pm}$, the expressions in (14) are in principle equivalent for all choices of reference. The freedom in $\Gamma_{0,T}$ and $\gamma^{\pm}$ allows us to ‘reweight’ between (14a), (14b) and (14c), or even cancel terms. A canonical choice is the Lebesgue measure for $\Gamma_{0}$ and $\Gamma_{T}$, and $\gamma^{\pm}=0$, see Appendix C.1.
@@
The distinction between forward and backward integration in (14) is related to the time points at which the integrands $\left(a_{t}-\gamma^{+}_{t}\right)({\bm{Y}}_{t})$ and $\left(b_{t}-\gamma^{-}_{t}\right)({\bm{Y}}_{t})$ would be evaluated in discrete-time approximations, e.g.,
$\int_{0}^{T}\!\!a_{t}({\bm{Y}}_{t})\!\cdot\!\overrightarrow{\mathrm{d}}{\bm{Y}}_{t}\approx\sum_{i}a_{t_{i}}({\bm{Y}}_{t_{i}})\!\cdot\!({\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}),\;\;\int_{0}^{T}\!\!a_{t}({\bm{Y}}_{t})\!\cdot\!\overleftarrow{\mathrm{d}}{\bm{Y}}_{t}\!\approx\!\sum_{i}a_{t_{i+1}}({\bm{Y}}_{t_{i+1}})\!\cdot\!({\bm{Y}}_{t_{i+1}}-{\bm{Y}}_{t_{i}}).$
Alternatively, forward and backward integrals can be transformed into each other using the conversion
@@
We refer to Kunita (2019) and Appendix A for further details. In passing, we note that (15) allows us to eliminate the Hutchinson estimator (Hutchinson, 1989)from a variety of common score-matching objectives, potentially reducing the variance of gradient estimators, see Appendix C.1.
@@
Framework 1 can be translated into the setting of (12), noting that (11) continues to hold with appropriate modifications:
Framework 1.
@@
For a divergence $D$ on path space, minimise $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})$. If $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})=0$, then (12a) transports $\mu$ to $\nu$, and (12b) transports $\nu$ to $\mu$.  [^3] Concurrently Richter & Berner (2024) propose an akin framework to ours.
@@
At optimality, $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})=0$, Proposition 2.1 allows us to obtain the scores associated to the learned diffusion via $\sigma^{2}\nabla\operatorname{ln}\rho_{t}^{\mu,a}=a_{t}-b_{t}$. In this way, Framework 1 is closely connected to (and in some ways extends) score-matching ideas (Song & Ermon, 2019; Song et al., 2021). Indeed, recent approaches towards generative modeling and sampling can be recovered from Framework 1 by making specific choices for the divergence $D$, the parameterisations for $a$ and $b$, as well as for the reference diffusion $\overrightarrow{{\mathbb{P}}}^{\Gamma_{0},\gamma^{+}}=\overleftarrow{{\mathbb{P}}}^{\Gamma_{T},\gamma^{-}}$ in Proposition 2.2:
@@
Score-based generative modeling: Letting $\mu$ be the target and fixing the forward drift $a_{t}$, and, motivated by Proposition 2.1, parameterising the backward drift as $b_{t}=a_{t}-s_{t}$, we recover the SGM objectives in Hyvärinen & Dayan (2005); Song & Ermon (2019); Song et al. (2021) from $D=D_{\mathrm{KL}}$; when $\overrightarrow{{\mathbb{P}}}^{\mu,a}=\overleftarrow{{\mathbb{P}}}^{\nu,b}$, the variable drift component $s_{t}$ will represent the score $\sigma^{2}\nabla\operatorname{ln}\rho^{\mu,a}_{t}$. Modifications can be obtained from the conversion formula (15), see Appendix C.2.
@@
Score-based sampling – ergodic drift: In this setting, $\nu$ becomes the target and we fix $b_{t}$ to be the drift of an ergodic (backward) process. Then choosing $\Gamma_{0,T}=\mu$, $\gamma^{\pm}=b$ allows us to recover the approaches in Vargas et al. (2023a); Berner et al. (2022). Possible generalisations based on Framework 1 include IWAE-type objectives, see Appendix C.3.
@@
Score-based sampling – Föllmer drift: Finally choosing $b_{t}(x)=x/t$ we recover Föllmer sampling (Appendix C.3; Follmer, 1984; Vargas et al., 2023b; Zhang & Chen, 2022; Huang et al., 2021b).
3 Learning forward and backward transitions simultaneously
Recall from the introduction that complete flexibility in $a$ and $b$ will render the minima of $D(\overrightarrow{{\mathbb{P}}}^{\mu,a}|\overleftarrow{{\mathbb{P}}}^{\nu,b})$ highly nonunique. Furthermore, the approaches surveyed at the end of the previous section circumvent this problem by fixing either $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ or $\overleftarrow{{\mathbb{P}}}^{\nu,b}$. However, to leverage the full power of diffusion models, both $\overrightarrow{{\mathbb{P}}}^{\mu,a}$ or $\overleftarrow{{\mathbb{P}}}^{\nu,b}$ should be adapted to the problem at hand. In this section, we explore models of this kind, by imposing additional constraints on $a$ and $b$. We end this section by presenting our new CMCD sampler connecting it to prior methodology within VI (Doucet et al., 2022b; Geffner & Domke, 2023; Papamakarios et al., 2017) and OT where we can view CMCD as an instance of entropy regularised OT in the infinite constraint limit (Bernton et al., 2019).
3.1 Connection to Entropic optimal transport
One way of selecting a particular transition between $\mu$ and $\nu$ is by imposing an entropic penalty, encouraging the dynamics to stay close to a prescribed, oftentimes physically or biologically motivated, reference process. Using the notation employed in Framework 1, the static Schrödinger problem (Schrödinger, 1931; Léonard, 2014a) is given by
$\pi^{*}({\bm{x}},{\bm{z}})\in\operatorname*{arg\,min}_{\pi({\bm{x}},{\bm{z}})}\Big{\{}D_{\mathrm{KL}}(\pi({\bm{x}},{\bm{z}})||r({\bm{x}},{\bm{z}})):\pi_{\bm{x}}({\bm{x}})=\mu({\bm{x}}),\pi_{\bm{z}}({\bm{z}})=\nu({\bm{z}})\Big{\}},$ (16)
where $r({\bm{x}},{\bm{z}})$ is the Schrödinger prior encoding additional domain-specific information. In an analogous way, we can introduce a regulariser to the path-space approach of Framework 1’ to obtain the dynamic Schrödinger problem